Calculate osmolality in the following example
The concentrations of solutes are:
[Na⁺]=0.150 mol/L;
[K⁺] = 0.005 mol/L
[Glucose]=380 mg/100 mL;
[BUN] (blood urea nitrogen)=28 mg/100 mL.
Q1: What is normal range
Q2. What are possibilities in this case
Answer #1

1.Calculated Serum osmolality formula is

Osmolality=2*Na + (Glucose/18) + (BUN/2.8)

In our case=2*150+(380/18)+(28/2.8)

=300+21.1+10

=331mOsm/Kg

2.Normal Range=285-295 mOsm/Kg

3.In this case Serum Osmolality is increased.Dehydration may be the cause of this. Hypernatremia is there.